# Optimizing Breast Cancer Detection in Mammograms: A Comprehensive Study of Transfer Learning, Resolution Reduction, and Multi-View Classification

### v1

@@ Abstract
This study explores open questions in the application of machine learning for breast cancer detection in mammograms. Current approaches often employ a two-stage transfer learning process: first, adapting a backbone model trained on natural images to develop a patch classifier, which is then used to create a single-view whole-image classifier. Additionally, many studies leverage both mammographic views to enhance model performance. In this work, we systematically investigate five key questions: (1) Is the intermediate patch classifier essential for optimal performance? (2) Do backbone models that excel in natural image classification consistently outperform others on mammograms? (3) When reducing mammogram resolution for GPU processing, does the “learn-to-resize” technique outperform conventional methods? (4) Does incorporating both mammographic views in a two-view classifier significantly improve detection accuracy? (5) How do these findings vary when analyzing low-quality versus high-quality mammograms? By addressing these questions, we developed models that outperform previous results for both single-view and two-view classifiers. Our findings provide insights into model architecture and transfer learning strategies contributing to more accurate and efficient mammogram analysis.
{IEEEkeywords}
Breast cancer detection, convolutional neural networks, deep learning, mammography, medical image analysis, multi-view classification, transfer learning.
@@ 1 Introduction
Breast cancer is the most common cancer among women globally, representing 23.76% of all new cancer cases diagnosed in 2022 [1]. Mammography, an X-ray-based imaging technique, is one of the most critical tools for early detection of the disease. To minimize error rates, mammograms should be interpreted by experienced radiologists, often supported by computer-aided detection and diagnosis (CAD) systems.
2 Literature review
CAD systems based on neural networks have been proposed for single-view mammogram analysis. For instance, Shen et al. [2] first train a patch-based classifier and then reuse its weights to train a full-image classifier. Shu et al. [3] introduced novel pooling techniques to enhance performance, while Wei et al. [4] proposed a morphing technique to resize mammograms for input into a ResNet [5] backbone with ImageNet pre-trained weights.
On the other hand, multi-view systems aim to improve performance by processing more than one view simultaneously. Approaches vary significantly: McKinney et al. [6] developed a system that outperformed radiologists using private datasets. Petrini et al. [7] proposed a two-view classifier that fuses CC (cranio-caudal) and MLO (mediolateral-oblique) features through 2D concatenation for convolutional processing, following the training of a patch-based and single-view classifier. Chen et al. [8] process global and local features from both views, extracted using a backbone network, for subsequent classification. Nguyen et al. [9] aggregate view features using operations like averaging and concatenation, followed by processing in a fully connected network. Sarker et al. [10] introduced a SwinTransformers-based approach that integrates self-attention and cross-attention mechanisms for breast mass lesion classification.
3 Methodology
@@
After analyzing multi-view mammogram classification approaches, we formulated the following key questions:
@@
    Is the patch classifier necessary? Shen et al. [2] and Petrini et al. [7] incorporated a patch classifier as an intermediate step in the transfer learning process. Is this step truly essential?
@@
By investigating these questions, we aim to refine best practices in machine learning-based mammogram classification, optimizing both accuracy and computational efficiency. In doing so, we developed models that surpass previous results for both single-view and two-view classifiers on the CBIS-DDSM dataset. To select the base model architectures, we will follow an evolutionary scale starting from ResNet [5] until ConvNeXt [14], using open source implementations from Wightman [15] and from PyTorch [^1] https://pytorch.org/vision/stable/models.html. We will use only publicly available mammogram datasets. Our source code and models are available online [^2] https://github.com/dpetrini/multiple-view.
4 Architecture Exploration in CBIS-DDSM
@@
4.1 Single View Classifiers
@@ 4.1.1 Patch Classifier and Base-Model
In this section, we aim to address questions (1) and (2) outlined in Section 3. To investigate these, we perform two sets of experiments using different approaches to initialize the weights of single-view classifiers:
@@ 4.1.3 Patch Based Classifier (PBC)
The resulting AUCs are presented in Table 3. Notably, we observed that the best-performing base models for patch classification did not consistently perform well for full mammogram classification. Specifically, while ConvNeXt-Base was the top-performing model for patch classification, EfficientNet-B3 emerged as the best base model for classifying whole single-view mammograms.
Table 3: AUCs obtained by single-view patch-based classifiers (PBC). PBC/DC indicates which approach generates the highest AUC in the best test, comparing with Table 4. “Train size” indicates the resolution of the training images of the base model.
@@
MobileNetV3-Large 224 0.7426$\pm$0.0065 0.7353$\pm$0.0205 DC
MobileNetV2 224 0.7504$\pm$0.0315 0.7611$\pm$0.0198 DC
ResNet-50 224 0.7541$\pm$0.0207 0.7616$\pm$0.0197 DC
ResNeXt-50-32x4d 224 0.7705$\pm$0.0068 0.7627$\pm$0.0197 PBC
DenseNet121 224 0.7804$\pm$0.0161 0.7633$\pm$0.0197 PBC
MNASNet-1.0 224 0.7627$\pm$0.0066 0.7643$\pm$0.0197 PBC
EfficientNet-B4 384 0.7866$\pm$0.0040 0.7814$\pm$0.0191 DC
ConvNeXt-Tiny 224 0.7969$\pm$0.0102 0.7850$\pm$0.0190 PBC
ResNet-18 224 0.7749$\pm$0.0129 0.7858$\pm$0.0189 PBC
EfficientNetV2-S 300 0.8034$\pm$0.0117 0.7886$\pm$0.0189 DC
DenseNet201 224 0.7884$\pm$0.0129 0.7891$\pm$0.0188 PBC
ResNet-101 224 0.7791$\pm$0.0093 0.7892$\pm$0.0188 PBC
EfficientNet-B0 224 0.7933$\pm$0.0096 0.7977$\pm$0.0185 DC
ConvNeXt-Small 224 0.8010$\pm$0.0033 0.7984$\pm$0.0185 PBC
ConvNeXt-Base 224 0.7991$\pm$0.0045 0.8019$\pm$0.0184 PBC
ConvNeXt-Base (22k) 224 0.8137$\pm$0.0108 0.8027$\pm$0.0183 PBC
EfficientNet-B2 260 0.8054$\pm$0.0128 0.8120$\pm$0.0180 PBC
DenseNet169 224 0.7910$\pm$0.0106 0.8031$\pm$0.0183 PBC
EfficientNet-B1 240 0.7884$\pm$0.0120 0.8054$\pm$0.0182 DC
EfficientNetV2-S (21k) 300 0.8089$\pm$0.0052 0.8114$\pm$0.0180 DC
EfficientNetV2-M 384 0.7931$\pm$0.0158 0.8150$\pm$0.0179 PBC
EfficientNet-B3 300 0.8114$\pm$0.0153 0.8325$\pm$0.0171 PBC
@@ 4.1.4 Direct Classifier (DC)
Table 4: AUCs obtained by single-view direct classifiers (DC). PBC/DC indicates which approach generates the highest AUC in the best test, comparing with Table 3. “Train size” indicates the resolution of the training images of the base model.
@@ 4.1.7 Comparison with Previous Work
The AUCs achieved by our patch-based classifier (PBC – 0.8325) and direct classifier (DC – 0.8313) are significantly higher than the best result reported in [7] (0.8033). Using the hypothesis test described on page 4.1.5, we compared the AUC of our PBC model with the previous work, obtaining a p-value of 0.0499. Similarly, the comparison for our DC model yielded a p-value of 0.0575. These results demonstrate that the improvements in our work are substantial, with some statistical significance. We attribute these gains to the use of modern pre-training strategies for ImageNet-based models. For instance, He et al. [18] improved the top-1 accuracy of ResNet50 from 75.3% to 79.29%, with 2.13% of this improvement directly attributable to advanced techniques such as cosine learning rate decay, label smoothing, and Mixup. Further enhancements by Wightman et al. [19] increased this accuracy to 80.4% through the incorporation of the LAMB optimizer and CutMix augmentation.
@@
4.2 Classifiers with Resized Images
To address question (3) outlined in Section 3, we tested reducing the input image size by half, resulting in dimensions of 576×448. We conducted two experiments using different downsampling methods. The first experiment employed conventional interpolation, while the second utilized a machine learning-based downsampling technique called “Learn to Resize” [11]. If this technique demonstrates superior performance compared to conventional resampling, it could become the preferred method for reducing high-resolution mammograms to a resolution compatible with current processing devices.
@@ 4.2.1 Fixed Resizing Classifier (FRC)
MNASNet-1.0 224 0.5643$\pm$0.0267 0.5860$\pm$0.0230
MobileNetV2 224 0.6858$\pm$0.0169 0.6972$\pm$0.0214
ConvNeXt-Small 224 0.7084$\pm$0.0240 0.7117$\pm$0.0211
ConvNeXt-Tiny 224 0.7118$\pm$0.0330 0.7214$\pm$0.0209
ResNet-18 224 0.7433$\pm$0.0274 0.7536$\pm$0.0200
EfficientNet-B2 260 0.7622$\pm$0.0143 0.7656$\pm$0.0196
EfficientNet-B1 240 0.7644$\pm$0.0077 0.7680$\pm$0.0195
DenseNet121 224 0.7896$\pm$0.0106 0.7751$\pm$0.0193
MobileNetV3-Large 224 0.7811$\pm$0.0131 0.7753$\pm$0.0193
EfficientNetV2-S (21k) 300 0.7808$\pm$0.0071 0.7761$\pm$0.0193
EfficientNetV2-M 384 0.7723$\pm$0.0269 0.7832$\pm$0.0190
ResNeXt-50-32x4d 224 0.7817$\pm$0.0137 0.7881$\pm$0.0189
EfficientNet-B4 384 0.7847$\pm$0.0053 0.7890$\pm$0.0188
DenseNet201 224 0.7813$\pm$0.0107 0.7943$\pm$0.0186
EfficientNet-B0 224 0.7686$\pm$0.0242 0.7956$\pm$0.0186
ConvNeXt-Base (22k) 224 0.7951$\pm$0.0057 0.7958$\pm$0.0186
DenseNet169 224 0.7795$\pm$0.0124 0.7968$\pm$0.0186
EfficientNet-B3 300 0.8010$\pm$0.0175 0.7972$\pm$0.0185
ResNet-101 224 0.7876$\pm$0.0207 0.7981$\pm$0.0185
ConvNeXt-Base 224 0.7423$\pm$0.0440 0.8009$\pm$0.0184
EfficientNetV2-S 300 0.7901$\pm$0.0102 0.8049$\pm$0.0183
ResNet-50 224 0.7756$\pm$0.0300 0.8167$\pm$0.0178

### v2

@@ Abstract
Mammography, an X-ray-based imaging technique, plays a crucial role in the early detection of breast cancer. Its accuracy heavily depends on expert radiologists, making it essential to minimize interpretation errors. To support radiologists, various computer-aided detection and diagnostic methods have been proposed, increasingly leveraging advancements in artificial intelligence and machine learning.
Over recent years, mammogram analysis has evolved significantly—from early patch-based classifiers, which examine only localized regions of images, to full-image classifiers, and later towards multi-view systems that simultaneously integrate multiple perspectives of the mammographic exam for enhanced accuracy. Despite this progression, critical questions remain, such as whether multi-view systems consistently outperform single-view approaches.
In this paper, we systematically evaluate and compare the effectiveness of single-view and multi-view mammogram classification techniques. Our research introduces models that achieve superior performance relative to existing state-of-the-art approaches in both single-view and two-view classification scenarios. Furthermore, our findings provide valuable insights into optimal model architectures and effective transfer learning strategies, paving the way for more accurate and efficient mammogram interpretation. The inference code and model are available at https://github.com/dpetrini/multiple-view.
@@ 1 Introduction
Breast cancer remains the most prevalent cancer among women globally, accounting for approximately 23.8% of all new cancer cases diagnosed in 2022 [1]. Mammography, an imaging technique based on X-rays, is currently the gold standard for early breast cancer detection. However, interpreting mammograms requires considerable expertise, and despite experienced radiologists utilizing computer-aided detection and diagnosis (CAD) systems, errors remain prevalent.
For instance, Bahl and Do [2] reported a sensitivity of only 87% in physician-interpreted mammograms, with an interval cancer rate—cancers diagnosed following a negative mammographic result—of approximately 30%. These figures highlight significant limitations in traditional diagnostic methods and emphasize the potential benefit of integrating AI-based CAD systems to reduce errors.
Initial research, such as the study by Kooi et al. [3], explored convolutional neural networks (CNNs) for classifying mammograms. However, direct application of CNN architectures designed for natural images did not initially yield promising results due to fundamental differences between natural images and mammograms. Specifically, natural images typically feature prominent objects occupying substantial portions of the image area, whereas cancerous lesions in mammograms are often subtle and localized to small regions. Moreover, natural images contain three color channels (Red, Green, Blue), whereas mammograms are single-channel grayscale images with high resolutions—commonly between 3,000 and 4,000 pixels in both dimensions. These high resolutions significantly increase computational complexity, challenging standard CNN-based classification techniques and conventional transfer learning strategies, thereby motivating the development of specialized methodologies.
To address these challenges, Shen et al. [4] proposed a two-stage transfer learning approach. Initially, they trained a "patch classifier" using ImageNet-pretrained CNNs [5] to categorize small mammogram regions (patches) into lesion types. Subsequently, this trained patch classifier initialized a "single-view whole-image classifier," which produced per-image cancer probability. Petrini et al. [6] expanded this approach by introducing a third stage that simultaneously analyzed both medio-lateral oblique (MLO) and cranio-caudal (CC) mammographic views, significantly improving accuracy when tested on public datasets.
Despite these advancements, it remains uncertain whether CNN architectures that achieve top performance on natural image benchmarks (e.g., ImageNet) will consistently outperform other models when applied to mammograms.
In this paper, we systematically explore several key questions influencing CNN-based mammogram classification: (1) training strategies (e.g., patch-based pretraining versus end-to-end learning), (2) choice of backbone architecture, (3) image resolution and downsampling techniques, (4) single-view versus multi-view integration, and (5) model performance on datasets with varying image quality.
Through addressing these questions, we propose enhanced models that achieve state-of-the-art performance for both single-view and two-view mammogram classification tasks. Our findings provide valuable insights into the choice of CNN architectures and transfer learning strategies, ultimately guiding the development of more effective and efficient mammographic analysis systems.
2 Materials & Methods
For breast cancer classification, machine learning studies may rely on proprietary mammography datasets, as in Wu et al. [7] and McKinney et al. [8], or on publicly accessible datasets. In this work, we exclusively utilize open-access mammographic repositories. In the next topic, we summarize the key characteristics of public datasets available, and in the following topics, we provide a literature review of techniques for breast cancer analysis, covering single view and multi-view architectures. We conclude this section describing the proposed methodology for enhancing mammogram classification performance.
2.1 Public Datasets
There are two principal types of mammography: screen-film mammography (SFM, conventional film-based) and full-field digital mammography (FFDM). To process SFM digitally, film images must first be scanned—further degrading their quality—whereas FFDM is inherently digital. A simple visual comparison of scanned SFM versus FFDM confirms that the latter offers substantially superior image quality. Accordingly, it is evident that training and evaluating an AI system on native digital mammograms will yield better results than using digitized analogue films.
Fair comparison of different computer-aided detection (CAD) systems for breast cancer requires large public FFDM datasets. However, early publicly available mammography repositories consist of low quality digitized SFM images: the Digital Database for Screening Mammography (DDSM) [9], the Mammographic Image Analysis Society database (MIAS) [10], and the Image Retrieval in Medical Applications project (IRMA) [11]. DDSM is the largest, comprising 2,620 exams with normal, benign, and malignant cases verified by pathology. The InBreast dataset [12] contains true FFDM images but includes only 115 cases (410 images), which is too small for deep-learning applications.
The CBIS-DDSM [13] is a curated and enhanced subset of the original DDSM, specifically tailored to support research applications. It comprises 3,103 mammographic images, benign or malignant, in this case confirmed by biopsy, and includes detailed lesion annotations to facilitate automated analysis.
The CSAW-M dataset [14], developed at the Karolinska Breast Center in Sweden, comprises 10,020 mammograms at 632×512-pixel resolution, annotated by at least one radiologist with a “masking” difficulty score rather than benign/malignant labels, reflecting lesions obscured by surrounding tissue.
The KAU-BCMD [15] from King Abdulaziz University in Saudi Arabia includes 1,416 exams (four views each, totaling 5,662 images) labeled by BI-RADS rather than gold-standard biopsy; 205 of these exams also include ultrasound images (BI-RADS 0), and the public download currently provides 2,378 mammograms at 1,267×578 resolution (version 10).
VinDr-Mammo [16] is a full-field digital mammography (FFDM) collection of 5,000 four-view exams (20,000 images) with BI-RADS assessments, breast density, and lesion annotations (masses, calcifications, asymmetries, architectural distortions); each exam was independently double-read with arbitration for discordance, and only BI-RADS 3–5 exams include lesion marks.
EMory BrEast Imaging Dataset (EMBED) [17] is another large scale FFDM dataset, comprises 3.65 million screening and diagnostic 2D and DBT mammograms, from racially diverse backgrounds. A portion of this dataset is available for download. It contains biopsy and ROIs (Region of Interest) pointing the lesions.
The RSNA dataset [18] was the subject of a recent Kaggle data challenge and contains 54,713 digital mammograms from almost 8,000 patients. It has indication of Bi-Rads, biopsy and has around 2,1% of positive cases.
The CMMD [19] Chinese Mammography Database contains 3,712 DICOM mammograms (2,294 × 1,914 pixels, 8-bit) from 1,775 patients, split into 2,214 images with biopsy-confirmed benign or malignant labels and 1,498 images with available molecular subtype information.
NL-Breast-Screening [20], is a recent Canadian population-screening FFDM dataset of 5,997 four-view exams, biopsy-confirmed, with positive, false-positive and normal labels but without lesions position indication, was released in 2024.
We selected the CBIS-DDSM, a conventional film-based dataset and the VinDr-Mammo, a full-field digital dataset, for the development of this work.
2.2 Related work
Below, we outline neural network–based CAD systems for single-view mammogram analysis to address the challenge of mammogram classification.
As mentioned earlier, Shen et al. [4] proposed an strategy based in two transfer-learning stages. They achieved an AUC of 0.87 using an specific selection of CBIS-DDSM images for test set. Other researchers [21] evaluated Shen’s method in the original test set and achieved an AUC of 0.75.
Shu et al. [22] proposed two novel pooling techniques to replace the conventional average-pooling and max-pooling layers, using the same dataset. Their method achieves a maximum AUC of 0.838. However, it is unclear whether they employed the original dataset split, since they report using an 85/15% train/test partition, whereas the original dataset is divided 79/21%.
Wei et al. [21], also using CBIS-DDSM dataset introduced the use of neural-network morphing in lieu of traditional transfer learning. Using the original train/test split, they report AUCs of 0.7964 for the single model without test-time augmentation (TTA), 0.8187 for the single model with TTA, and 0.8313 for an ensemble of four models with TTA. By contrast, they obtain an AUC of 0.9427 when employing a fixed random train/test split; these results are not directly comparable due to different test splits.
Panceri et al. [23] conducted a study in which they used only 503 craniocaudal (CC) views containing calcification lesions, randomly selected from the 717 CBIS-DDSM images meeting these criteria. Their approach is patch-based: they extract 256 × 256 patches from the selected images, trained a patch classifier, and then apply this classifier to entire test images by dividing them into patches and classifying each one. They aggregate the per-patch probabilities to produce an overall mammography diagnosis.
Bhat et al. [24] employ a technique called AUCReshaping, which aims to modify the ROC curve to improve evaluation within specific sensitivity and specificity ranges. Using the VinDr-Mammo dataset, they first classify breast density and report a high AUC of 0.97. They also evaluate BI-RADS categories, grouping 1, 2, and 3 as the negative class and 4 and 5 as the positive class, achieving an AUC of 0.77. This grouping yields a different class balance than our further experiments.
On the other hand, multi-view systems aim to improve performance by processing more than one view simultaneously. Approaches vary significantly as below.
Khan et al. [25] introduce a multi-view, patch-based approach using a simplified VGG-style convolutional network to analyze mammographic ROIs from the MIAS and CBIS-DDSM datasets. In three sequential training stages, their model first distinguishes normal from abnormal patches on MIAS with an AUC of 0.934, then classifies CBIS-DDSM patches as masses versus calcifications, AUC of 0.923, and finally differentiates benign from malignant lesions, AUC of 0.769.
Petrini et al. [6] enhanced the results of Shen et al. work [4] by adopting a more advanced convolutional backbone and introducing a two-view classifier. In this design, CC (craniocaudal) and MLO (mediolateral-oblique) feature maps are fused via 2D concatenation and processed jointly by convolutional layers, following the pretraining of a patch-based model and a single-view classifier. Evaluated on the CBIS-DDSM dataset using its original train/test split, the two-view system achieved an AUC of 0.8418±0.0258.
Chen et al. [26] proposed a multi-view system that uses an EfficientNet-b0 backbone, resized images to 1,536×768, and processed independently global features and region-level features, passing the later through multi-head attention, from both CC and MLO views, then fuses both for subsequent classification. They trained in a private dataset and tested in CMMD [19] (complete dataset) and InBreast [12] (test division).
Nguyen et al. [27] introduced a two-view classification framework that feeds paired CC and MLO mammograms into parallel ResNet backbones. The extracted feature maps are merged in a fusion block—via averaging, concatenation, and element-wise addition—and subsequently classified by fully connected layers. When evaluated on the CMMD and VinDr-Mammo datasets (with BI-RADS 2 as the negative class and BI-RADS 5 as the positive class), this approach achieved a peak AUC of 0.7486.
Sarker et al. [28] introduced a Swin Transformer Multi-Headed Dynamic Attention module, resized all images to 224×224 and 384×384 and evaluated on CBIS-DDSM and VinDr-Mammo datasets. This approach, exclusively based on transformers, achieved an AUC of 71.37 using the two views mass mammograms from CBIS-DDSM.
2.3 Methodology
@@
In this research paper, after analyzing multi-view mammogram classification approaches, we formulated the following key questions:
@@
    Are patch classifiers necessary? Some approaches like Shen et al. [4] and Petrini et al. [6] incorporated a patch classifier as an intermediate step in the transfer learning process. Is this step truly essential?
@@
By investigating these questions, we aim to refine best practices in machine learning-based mammogram classification, optimizing both accuracy and computational efficiency. In doing so, we developed models that surpass previous results for both single-view and two-view classifiers on the CBIS-DDSM and VinDr-Mammo dataset. To select the base model architectures, inspired by [30], we will follow an evolutionary scale starting from ResNet [31] until ConvNeXt [32], using open source implementations from Wightman [33] and from PyTorch [34] and will evaluate different training approaches, like patch-based-pre-train, resizing and others, as summarized in Figure 1.
Figure 1: Exploration methodology.
We will use only publicly available mammogram datasets. Our source code and models are available online [35].
3 Results and Discussion
In this section we apply the methodology in both public datasets, firstly in CBIS-DDSM and later in VinDr-Mammo.
3.1 Architecture Exploration in CBIS-DDSM
@@
3.2 Single View Classifiers in CBIS-DDSM
@@ 3.2.1 Patch Classifier and Base-Model
In this section, we aim to address questions (1) and (2). To investigate these, we perform two sets of experiments using different approaches to initialize the weights of single-view classifiers:
@@ 3.2.3 Patch Based Classifier (PBC)
The resulting AUCs are presented in Table 3. Notably, we observed that the best-performing base models for patch classification did not consistently perform well for full mammogram classification. Specifically, while ConvNeXt-Base was the top-performing model for patch classification, EfficientNet-B3 emerged as the best base model for classifying whole single-view mammograms. In this table and in following ones, “Train size” indicates the resolution of the pre-training images of the base model.
Table 3: AUCs obtained by single-view patch-based classifiers (PBC). PBC/DC indicates which approach generates the highest AUC in the best test, comparing with Table 4.
@@
[…]
@@ 3.2.4 Direct Classifier (DC)
Table 4: AUCs obtained by single-view direct classifiers (DC). PBC/DC indicates which approach generates the highest AUC in the best test, comparing with Table 3.
@@ 3.2.7 Comparison with Previous Work
The AUCs achieved by our patch-based classifier (PBC – 0.8325) and direct classifier (DC – 0.8313) are significantly higher than the best result reported in [6] (0.8033). Using the hypothesis test described on page 3.2.5, we compared the AUC of our PBC model with the previous work, obtaining a p-value of 0.0499. Similarly, the comparison for our DC model yielded a p-value of 0.0575. These results demonstrate that the improvements in our work are substantial, with some statistical significance. We attribute these gains to the use of modern pre-training strategies for ImageNet-based models. For instance, He et al. [38] improved the top-1 accuracy of ResNet50 from 75.3% to 79.29%, with 2.13% of this improvement directly attributable to advanced techniques such as cosine learning rate decay, label smoothing, and Mixup. Further enhancements by Wightman et al. [39] increased this accuracy to 80.4% through the incorporation of the LAMB optimizer and CutMix augmentation.
@@
3.3 Classifiers with Resized in CBIS-DDSM
To address question (3), we tested reducing the input image size by half, resulting in dimensions of 576×448. We conducted two experiments using different downsampling methods. The first experiment employed conventional interpolation, while the second utilized a machine learning-based downsampling technique called “Learn to Resize” [29]. If this technique demonstrates superior performance compared to conventional resampling, it could become the preferred method for reducing high-resolution mammograms to a resolution compatible with current processing devices.
@@ 3.3.1 Fixed Resizing Classifier (FRC)
[…]